# Susceptibilities of rotating quark matter in Fourier-Bessel basis

### v1

@@ IV.1 Meson and topological susceptibilities
Besides, we evaluate the topological susceptibility under rotation by the analytical expression in Eq. (55) with the numerical results of the pseudoscalar meson susceptibilities obtained above. Fig. 4 presents the $r$-dependence in panel (a) and the temperature dependence in panel (b). These dependences closely resemble the behaviors of the pseudoscalar meson susceptibilities shown in Fig. 3. From these results, the role of the rotational effect on the topological susceptibility can be understood as follows: it further facilitates the thermal suppression of the topological susceptibility and becomes more pronounced near the boundary, following a trend similar to that of the pseudoscalar meson susceptibilities.
In addition, the behavior of the topological susceptibility in a rotating system can also be understood as follows. As shown in Fig. 3, the $\eta$ meson susceptibility is relatively small, whereas the pion susceptibility is significantly larger, leading to $\chi_{\pi}$ being the dominant contribution to the topological susceptibility. Notably, the pion susceptibility can be expressed in terms of the quark condensate via the WT identity in Eq. (30). Therefore, the behavior of the topological susceptibility reflects that of the quark condensate under rotation. Indeed, a similar correspondence between the behavior of the topological susceptibility and that of the quark condensate has also been observed in other contexts, such as finite temperature systems [67, 68, 69, 50, 45, 46, 70] and finite-density systems [71, 47].
@@
IV.2 Baryon number susceptibility
@@
In Fig. 6, we numerically compute the baryon number susceptibility. To investigate the influence of the chiral phase transition, we compare the susceptibilities in the interacting and noninteracting cases. The former (latter) is denoted by $\chi_{B}^{{\rm w}/M}$ ($\chi_{B}^{{\rm w}/m_{0}}$) and plotted with the solid (dashed) lines; the latter is evaluated by replacing the dynamical quark mass $M$ in the Fermi distribution function $n_{F}(E)$ of Eq. (57) with the current quark mass $m_{0}$. In the absence of the rotational effect, while for $T\lesssim T_{\rm pc}$, there are discrepancies between the two cases, as depicted in panel (a), such discrepancies vanish for $T\gtrsim T_{\rm pc}$. This is because the dynamical quark mass approaches the current quark mass at high temperatures. To clearly capture this trend, in panel (b) we plot $\chi_{B}^{{\rm w}/M}/\chi_{B}^{{\rm w}/m_{0}}$, which is the susceptibility normalized by that in the noninteracting case. When including the rotational effects, this convergence shifts to lower-temperature regions, as shown in panel (b). The baryon number susceptibility is correlated with the chiral restoration affected by the rotational effect.
@@
IV.3 Moment of inertia
In Fig. 7, we show numerical results of the moment of inertia (58), which is independent of $\Omega$. Panel (a) is the radial-coordinate dependence, and the blue and red lines correspond to the symmetry-broken and symmetry-restored cases, respectively. First, we focus on the latter. While the moment of inertia is tiny around $r=0$, it grows for larger $r$ and reaches a peak around $r=0.9\,R$. This behavior stems from the contributions from the modes that has a large total angular momentum $j=l+1/2$, and thus spatially locate around a large $r$. Such a different location of each mode is the reason why the moment of inertia exhibits the clear $r$-dependence even at $\Omega=0$, while other susceptibility functions are insensitive to $r$ at $\Omega=0$. On the other hand, the blue line in panel (a) exhibits the small value irrelevantly to $r$, because the fermionic distribution is suppressed by a huge dynamical mass. This observation is supported by panel (b), where the moment of inertia is a monotonically increasing function of $T$. We note that the difference between the green and yellow lines again stems from the contribution from large $j$-modes.
@@ V Discussion and summary
Although we have adopted the mean field approximation supplemented by the local density approximation, this study presents the first analysis of the susceptibility functions of rotating quark fields, properly taking into account the violation of the translational invariance. The susceptibility functions are measurable in lattice QCD simulations even under finite angular velocity, employing techniques such as the imaginary angular velocity method [35] and Taylor expansion analysis [38]. Furthermore, the moment of inertia evaluated at zero angular velocity is not only a computable quantity without suffering from a sign problem, but also an essential ingredient characterizing the rotational response on the vacuum [36, 21]. An important remark is that moment of inertia is strongly affected by the geometry of systems, as well as boundary conditions. For instance, the orbital contribution is always washed out in the Cartesian coordinate system with the spatially periodic boundary condition, unlike our analysis in this paper. Including the treatment of such an inhomogeneity in lattice gauge theory, more comprehensive discussions could bridge a gap between lattice simulations of rotating QCD matter and other approaches.
Acknowledgements.

### v2

@@ 4.1 Meson and topological susceptibilities
Besides, we evaluate the topological susceptibility under rotation by the analytical expression in Eq. (3.31) with the numerical results of the pseudoscalar meson susceptibilities obtained above. Fig. 4 presents the $r$-dependence in panel (a) and the temperature dependence in panel (b). These dependences closely resemble the behaviors of the pseudoscalar meson susceptibilities shown in Fig. 3. From these results, the role of the rotational effect on the topological susceptibility can be understood as follows: it further facilitates the thermal suppression of the topological susceptibility and becomes more pronounced near the boundary, following a trend similar to that of the pseudoscalar meson susceptibilities.
In addition, the behavior of the topological susceptibility in a rotating system can also be understood as follows. As shown in Fig. 3, the $\eta$ meson susceptibility is relatively small, whereas the pion susceptibility is significantly larger, leading to $\chi_{\pi}$ being the dominant contribution to the topological susceptibility. Notably, the pion susceptibility can be expressed in terms of the quark condensate via the WT identity in Eq. (3.7). Therefore, the behavior of the topological susceptibility reflects that of the quark condensate under rotation. Indeed, a similar correspondence between the behavior of the topological susceptibility and that of the quark condensate has also been observed in other contexts, such as finite temperature systems Ruivo:2011fg; Jiang:2012wm; GomezNicola:2019myi; Kawaguchi:2020kdl; Kawaguchi:2020qvg; Cui:2021bqf; Cui:2022vsr; Fejos:2025oxi (for details of functional renormalization group analysis, see also Ref. Fejos:2016hbp) and finite-density systems Li:2022dry; Kawaguchi:2023olk.
@@
4.2 Baryon number susceptibility
@@
In Fig. 6, we numerically compute the baryon number susceptibility. To investigate the influence of the chiral phase transition, we compare the susceptibilities in the interacting and noninteracting cases. The former (latter) is denoted by $\chi_{B}^{{\rm w}/M}$ ($\chi_{B}^{{\rm w}/m_{0}}$) and plotted with the solid (dashed) lines; the latter is evaluated by replacing the dynamical quark mass $M$ in the Fermi distribution function $n_{F}(E)$ of Eq. (3.33) with the current quark mass $m_{0}$. In the absence of the rotational effect, while for $T\lesssim T_{\rm pc}$, there are discrepancies between the two cases, as depicted in panel (a), such discrepancies vanish for $T\gtrsim T_{\rm pc}$. This is because the dynamical quark mass approaches the current quark mass at high temperatures. To clearly capture this trend, in panel (b) we plot $\chi_{B}^{{\rm w}/M}/\chi_{B}^{{\rm w}/m_{0}}$, which is the susceptibility normalized by that in the noninteracting case. When including the rotational effects, this convergence shifts to lower-temperature regions, as shown in panel (b). The baryon number susceptibility is correlated with the chiral restoration affected by the rotational effect.
@@
4.3 Moment of inertia
In Fig. 7, we show numerical results of the moment of inertia (3.34), which is independent of $\Omega$. Panel (a) is the radial-coordinate dependence, and the blue and red lines correspond to the symmetry-broken and symmetry-restored cases, respectively. First, we focus on the latter. While the moment of inertia is tiny around $r=0$, it grows for larger $r$ and reaches a peak around r=0.9Rr=0.9\,R. This behavior stems from the contributions from the modes that has a large total angular momentum $j=l+1/2$, and thus spatially locate around a large $r$. Such a different location of each mode is the reason why the moment of inertia exhibits the clear $r$-dependence even at $\Omega=0$, while other susceptibility functions are insensitive to $r$ at $\Omega=0$. On the other hand, the blue line in panel (a) exhibits the small value irrelevantly to $r$, because the fermionic distribution is suppressed by a huge dynamical mass. This observation is supported by panel (b), where the moment of inertia is a monotonically increasing function of $T$. We note that the difference between the green and yellow lines again stems from the contribution from large $j$-modes.
@@ 5 Discussion and summary
As a side remark, we note that a recent study has computed susceptibilities in the context of investigating the critical end point in the QCD phase diagram Chen:2024hki. The authors evaluate the quark–antiquark spin correlation, which is formulated in terms of quantities analogous to the baryon number susceptibility and the moment of inertia. In their approach, separate chemical potentials and angular velocities are introduced for quarks and antiquarks, from which they derive corresponding susceptibility functions. As a result of this independent treatment of quarks and antiquarks, the physical quantities they obtain differ from the baryon number susceptibility and moment of inertia considered in our analysis. Besides, the susceptibility functions evaluated in this paper is the local one, unlike that in Ref. Chen:2024hki.
Although we have adopted the mean field approximation supplemented by the local density approximation, this study presents the first analysis of the susceptibility functions of rotating quark fields, properly taking into account the violation of the translational invariance. The susceptibility functions are measurable in lattice QCD simulations even under finite angular velocity, employing techniques such as the imaginary angular velocity method Braguta:2022str and Taylor expansion analysis Yang:2023vsw. Furthermore, the moment of inertia evaluated at zero angular velocity is not only a computable quantity without suffering from a sign problem, but also an essential ingredient characterizing the rotational response on the vacuum Braguta:2023yjn; Mameda:2023sst. An important remark is that moment of inertia is strongly affected by the geometry of systems, as well as boundary conditions. For instance, the orbital contribution is always washed out in the Cartesian coordinate system with the spatially periodic boundary condition, unlike our analysis in this paper. Including the treatment of such an inhomogeneity in lattice gauge theory, more comprehensive discussions could bridge a gap between lattice simulations of rotating QCD matter and other approaches.
Acknowledgment